OpenStudy (anonymous):

lim_(x->0^+) (x^2/2-1/x) How do I figure out this problem? Find the limit.

OpenStudy (anonymous):

\[\frac{ x^2 }{ 2-1/x } = \frac{ x^3 }{ 2x-1 }\]

OpenStudy (anonymous):

i didn't get your question.

OpenStudy (anonymous):